A small regional carrier accepted 14 reservations for a particular flight with 12 seats. 6 reservations went to regular customers who will arrive for the flight. Each of the remaining passengers will arrive for the flight with a 59% chance, independently of each other. (Report answers accurate to 4 decimal places.)

Find the probability that overbooking occurs.

Find the probability that the flight has empty seats.
Answer #1

Ans:

Out of 14 reservations,6 will arrive for sure.

Now,rest 8 passengers have probability of showing up,p=0.53

so use binomial distribution with n=10 and p=0.59

(out of 12 seats,6 will be occupied,so,x=6)

P(overbooked)=P(X>6)=1-P(X<=6)

=1-binomcdf(8,0.59,6)

=1-0.90368 =0.09632

P(empty seats)=P(X<6)=P(X<=5)=0.7052